# Approximate symmetries, insulators, and superconductivity in continuum-model description of twisted WSe2

### v1

@@ Abstract
Motivated by the recent discovery of superconductivity in twisted bilayer WSe2, we analyze the correlated physics in this system in the framework of a continuum model for the moiré superlattice. Using the symmetries in a fine-tuned limit of the system, we identify the strong-coupling ground states and their fate when the perturbations caused by finite bandwidth, displacement field, and the phase of the intralayer potential are taken into account. We classify the superconducting instabilities and, employing a spin-fermion-like model, study the superconducting instabilities in proximity to these insulating particle-hole orders. This reveals that only a neighboring intervalley coherent phase (with zero or finite wave vector) is naturally consistent with the observed superconducting state. Depending on details, the superconductor will be nodal or a chiral gapped state while further including electron-phonon coupling leads to a fully gapped, time-reversal symmetric pairing state.
@@
Inspired by the strong-coupling analysis above, we start by assuming that the (static) susceptibility $\chi(\boldsymbol{q})$ of the collective bosons is peaked at $\boldsymbol{q}=0$. The boson $\phi_{z}$ will mediate a repulsive intra-valley Cooper channel interaction and is, thus, not expected to stabilize superconductivity. To demonstrate that, we solve the linearized gap equations (LGE) numerically, using $\chi(\boldsymbol{q})=C\frac{\alpha}{\alpha^{2}+\boldsymbol{q}^{2}/k_{\theta}^{2}}$ and taking $\lambda_{j}$ in Eq. (7) to be the projections of $\eta_{j}$ onto the Bloch states of the band at the Fermi level See Appendix for more information ; here $\alpha$ parametrizes the bosonic mass which becomes smaller when approaching the neighboring insulating phase and we take $\alpha=1$ and $C/A_{\text{moire}}=1.7$ meV unless otherwise specified. Indeed, we find vanishingly small eigenvalues for the case of fluctuations of $\phi_{z}$. This is to be contrasted with fluctuations of $\phi_{x,y}$, which mediate an attractive intervalley interaction, in the sense that they favor $\Delta_{\boldsymbol{k},\eta}\simeq\Delta_{\boldsymbol{k},-\eta}$. However, due to the Fermi-Dirac constraint $\Delta_{\boldsymbol{k},\eta}=-\Delta_{-\boldsymbol{k},-\eta}$ and the significantly larger band separation compared to twisted bi- and trilayer graphene, where interband pairing is stabilized by these types of interactions Christos et al. (2023), we here have $\Delta_{\boldsymbol{k}}\simeq-\Delta_{-\boldsymbol{k}}$ and, thus, sign changes of the order parameter in the Brillouin zone. While this indicates that the order parameter transforms under an $\mathcal{I}$-odd representation for $D\rightarrow 0$, all of $A_{1}^{-}$, $A_{2}^{-}$, and $E^{-}$ are in principle possible and the outcome depends on energetics. In line with these considerations, we find in our numerics in Fig. 2(a,b) that the states with largest pairing strength (eigenvalue) not only emanate from $\mathcal{I}$-odd IRs [see (a)] but also stay approximately odd in $\boldsymbol{k}$ in the experimentally relevant range of $D$, see Fig. 2(c-e). We further observe that the $A_{1}^{-}$ and $E^{-}$ states are competing and which of the two is favored depends on the proximity to the insulator (the value of $\alpha$). As a result of the additional symmetry-imposed lines of zeros of $\Delta_{\boldsymbol{k}}$ at the zone boundary, the $A_{2}^{-}$ state in Fig. 2(e) has very small eigenvalues.
@@ Acknowledgements.
M.S.S. acknowledge funding by the European Union (ERC-2021-STG, Project 101040651—SuperCorr). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. M.C. and P.M.B. acknowledge funding from U.S. National Science Foundation grant No. DMR-2245246. M.C. thanks Ming Xie for helpful conversations on pairing mechanisms.

### v2

@@ Abstract
Motivated by the recent discovery of superconductivity in twisted bilayer WSe2, we analyze the correlated physics in this system in the framework of a continuum model for the moiré superlattice. Using the symmetries in a fine-tuned limit of the system, we identify the strong-coupling ground states and their fate when the perturbations caused by finite bandwidth, displacement field, and the phase of the intralayer potential are taken into account. We classify the superconducting instabilities and, employing a spin-fermion-like model, study the superconducting instabilities in proximity to these insulating particle-hole orders. This reveals that only a neighboring intervalley coherent phase (with zero or finite wave vector) is naturally consistent with the observed superconducting state, which we show to be crucially affected by the non-trivial band topology. Depending on details, the superconductor will be nodal or a chiral gapped state while further including electron-phonon coupling leads to a fully gapped, time-reversal symmetric pairing state.
@@
Inspired by the strong-coupling analysis above, we start by assuming that the (static) susceptibility $\chi(\boldsymbol{q})$ of the collective bosons is peaked at $\boldsymbol{q}=0$. The boson $\phi_{z}$ will mediate a repulsive intra-valley Cooper channel interaction and is, thus, not expected to stabilize superconductivity. To demonstrate that, we solve the linearized gap equations (LGE) numerically, using $\chi(\boldsymbol{q})=C\frac{\alpha}{\alpha^{2}+\boldsymbol{q}^{2}/k_{\theta}^{2}}$ and taking $\lambda_{j}$ in Eq. (7) to be the projections of $\eta_{j}$ onto the Bloch states of the band at the Fermi level See Appendix for more information ; here $\alpha$ parametrizes the bosonic mass which becomes smaller when approaching the neighboring insulating phase and we take $\alpha=1$ and $C/A_{\text{moire}}=1.7$ meV unless otherwise specified. Indeed, we find vanishingly small eigenvalues for the case of fluctuations of $\phi_{z}$. This is to be contrasted with fluctuations of $\phi_{x,y}$, which mediate an attractive intervalley interaction, in the sense that they favor $\Delta_{\boldsymbol{k},\eta}\simeq\Delta_{\boldsymbol{k},-\eta}$. However, due to the Fermi-Dirac constraint $\Delta_{\boldsymbol{k},\eta}=-\Delta_{-\boldsymbol{k},-\eta}$ and the significantly larger band separation compared to twisted bi- and trilayer graphene, where interband pairing is stabilized by these types of interactions Christos et al. (2023), we here have $\Delta_{\boldsymbol{k}}\simeq-\Delta_{-\boldsymbol{k}}$ and, thus, sign changes of the order parameter in the Brillouin zone. While this indicates that the order parameter transforms under an $\mathcal{I}$-odd representation for $D\rightarrow 0$, all of $A_{1}^{-}$, $A_{2}^{-}$, and $E^{-}$ are in principle possible and the outcome depends on energetics. In line with these considerations, we find in our numerics in Fig. 2(a,b) that the states with largest pairing strength (eigenvalue) not only emanate from $\mathcal{I}$-odd IRs [see (a)] but also stay approximately odd in $\boldsymbol{k}$ in the experimentally relevant range of $D$, see Fig. 2(c-e). We further observe that the $A_{1}^{-}$ and $E^{-}$ states are competing and which of the two is favored depends on the proximity to the insulator (the value of $\alpha$). We note that the dominance of $A_{1}$ pairing for larger $\alpha$ is crucially related to the non-zero Chern number of the bands, which requires $\lambda_{j}(\boldsymbol{k},\boldsymbol{q})$ to vanish at $\boldsymbol{k}=\boldsymbol{q}=0$ See Appendix for more information . As a result of the additional symmetry-imposed lines of zeros of $\Delta_{\boldsymbol{k}}$ at the zone boundary, the $A_{2}^{-}$ state in Fig. 2(e) has very small eigenvalues.
@@ Acknowledgements.
M.S.S. acknowledge funding by the European Union (ERC-2021-STG, Project 101040651—SuperCorr). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. M.C. and P.M.B. acknowledge funding from U.S. National Science Foundation grant No. DMR-2245246. P.M.B. acknowledges support by the German National Academy of Sciences Leopoldina through Grant No. LPDS 2023-06. M.C. thanks Ming Xie for helpful conversations on pairing mechanisms.
@@ Appendix B Additional numerical results
Appendix C Topological constraints on form factors and impact on pairing
In this appendix, we discuss how the non-zero Chern number of the active bands necessarily leads to a non-trivial momentum dependence of the form factors. As IVC fluctuations provide the most promising pairing glue in our theory, we illustrate this using the intervalley form factors $\lambda^{+-}_{x,y}(\boldsymbol{k},\boldsymbol{q})$ entering the associated LGE in Eq. (11). Using time-reversal symmetry to relate the wave functions in the two valleys, the form factors can be written as $\lambda^{+-}_{x}(\boldsymbol{k},\boldsymbol{q})=\braket{u^{*}_{+}(-\boldsymbol{k})}{u_{+}(\boldsymbol{k}+\boldsymbol{q})}$, and similarly for the $y$ component, where |$\ket{u_{\eta}(\boldsymbol{k})}$ are the Bloch states in valley $\eta$. Viewing $\lambda^{+-}_{x}(\boldsymbol{k},\boldsymbol{q})$ as a function $s_{\boldsymbol{k}_{0}}(\boldsymbol{q})=\lambda^{+-}_{x}(\boldsymbol{k}_{0},\boldsymbol{q})$ of the momentum transfer $\boldsymbol{q}$ at fixed $\boldsymbol{k}=\boldsymbol{k}_{0}$, we conclude that $s_{\boldsymbol{k}_{0}}(\boldsymbol{q})$ has to vanish at least for one $\boldsymbol{q}$ in the Brillouin zone. Since, otherwise, we could use |$\ket{u^{*}_{+}(-\boldsymbol{k}_{0})}$ as a (momentum $\boldsymbol{q}$ independent) “trial Bloch state” that, upon projection to the band $\{u_{+}(\boldsymbol{k}_{0}+\boldsymbol{q})\}_{\boldsymbol{q}}$, would define a smooth and periodic gauge for that band. This would be at odds with the non-zero Chern number PhysRevB.74.235111.
Three-fold rotational symmetry further implies that $s_{\boldsymbol{k}_{0}}(\boldsymbol{q})=s_{C_{3z}\boldsymbol{k}_{0}}(C_{3z}\boldsymbol{q})$. Therefore, for high-symmetry choices of $\boldsymbol{k}_{0}$ with $C_{3z}\boldsymbol{k}_{0}=\boldsymbol{k}_{0}$, these zeros are either pinned at high-symmetry $\boldsymbol{q}$, which are themselves $C_{3z}$ invariant, or appear in multiples of three. These expectations are confirmed by Fig. 10, where the intervalley form factors for the active band at the Fermi surface are shown as a function of $\boldsymbol{q}$ at fixed $\boldsymbol{k}$. For the high-symmetry choices $\boldsymbol{k}=\Gamma$ and $K$, the zeros are indeed pinned at high-symmetry points, too. While generic $\boldsymbol{k}$ leads to zeros at generic positions.
Based on the LGE in Eq. (11), it is natural that these topology-related momentum dependencies in the form factors influence the superconducting energetics. From Fig. 2(b), we can see that making the peak of $\chi_{\boldsymbol{q}}$ more concentrated around $\boldsymbol{q}=0$ by decreasing $\alpha$ favors the $E$ state. As such, it seems natural that the zero at $\boldsymbol{q}=0$ in Fig. 10(a) is an important driving force for the $A_{1}$ state’s dominance in Fig. 2. To test this hypothesis, we have removed the momentum-dependence of the form factors, $\lambda_{x,y}\rightarrow 1$, in the LGE and, indeed, this favors the $E$ pairing instability, see Fig. 11.
Figure 10: We show the quantity $\chi_{\boldsymbol{q}}|\langle u_{+}(\boldsymbol{k})|u_{-}(\boldsymbol{k}+\boldsymbol{q})\rangle|^{2}$ as a function of $\boldsymbol{q}$ in the first Brillouin zone for three different values of $\boldsymbol{k}$. In (a) we choose $\boldsymbol{k}=\Gamma$, in (b) $\boldsymbol{k}$ to lies halfway between $\Gamma$ and $K$, and in (c) $\boldsymbol{k}=K$. We use $\chi_{\boldsymbol{q}}$ peaked at $\boldsymbol{q}=0$ as in the main text and $D/d=14.45$ mV/nm.
Figure 11: Solutions of the LGE for fluctuations of $\phi_{x,y}$ when we set the form factors $\lambda_{x,y}(\boldsymbol{k},\boldsymbol{q})=1$ to study the impact of the strong momentum dependence of $\lambda_{x,y}$ demonstrated in Fig. 10. As in Fig. 7, (a) shows the leading eigenvalues for each IR, while (b) and (c) show the corresponding order parameters in the Brillouin zone for the leading and subleading solution, respectively. Comparison with Fig. 2(b) of the main text shows that removing the momentum dependence of the form factors favors the $E$ state relative to the $A_{1}$ state.